### v1

# TALE: A Tool-Augmented Framework for Reference-Free Evaluation of Large Language Models

Abstract
As Large Language Models (LLMs) become increasingly integrated into real-world, autonomous applications, relying on static, pre-annotated references for evaluation poses significant challenges in cost, scalability, and completeness. We propose Tool-Augmented LLM Evaluation (TALE), a framework to assess LLM outputs without predetermined ground-truth answers. Unlike conventional metrics that compare to fixed references or depend solely on LLM-as-a-judge knowledge, TALE employs an agent with tool-access capabilities that actively retrieves and synthesizes external evidence. It iteratively generates web queries, collects information, summarizes findings, and refines subsequent searches through reflection. By shifting away from static references, TALE aligns with free-form question-answering tasks common in real-world scenarios. Experimental results on multiple free-form QA benchmarks show that TALE not only outperforms standard reference-based metrics for measuring response accuracy but also achieves substantial to near-perfect agreement with human evaluations. TALE enhances the reliability of LLM evaluations in real-world, dynamic scenarios without relying on static references.
1 Introduction
Recent progress in Large Language Models (LLMs) has led to systems capable of producing fluent, context-aware, and semantically rich text across a wide range of tasks. Yet, as these models grow in complexity and capability, achieving timely, affordable, and accurate evaluation of their outputs remains a significant challenge (Xu et al., 2025).
Consider the question: “What is the tallest building in the world?” The correct answer has changed over time—from the “Willis Tower” to “Taipei 101”, and more recently, to the “Burj Khalifa.” and could be expressed in various valid ways: “Burj Khalifa,” “Khalifa Tower,” “Burj Dubai,” or with its precise height. Traditional reference-based evaluation metrics, including Exact Match (EM), F1, BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004), rely on fixed ground-truth references and thus may incorrectly penalize such legitimate variations (Kamalloo et al., 2023; Wang et al., 2023). These methods, though efficient, struggle to reflect semantic equivalence, are limited in adaptability to evolving facts. On the other hand, human evaluation is generally reliable, but it becomes overly expensive and labor-intensive when applied at scale (Chiang & Lee, 2023; Mañas et al., 2024; Zhu et al., 2023).
The challenge of evaluating LLMs grows as they act as autonomous agents, handling tasks like web browsing, planning, and information synthesis (Xi et al., 2023; Fu et al., 2023). In such dynamic settings, outputs are often unpredictable, context-dependent, and non-deterministic, making it impractical to pre-annotate reference answers for every possible interaction (Li et al., 2024). As a result, static, reference-driven evaluation protocols are fundamentally misaligned with the needs of modern LLM applications operating in open-ended, real-world environments.
An emerging alternative to static, reference-based evaluation is the LLM-as-a-judge approach (Zheng et al., 2024), where one model, for instance, is prompted to assess the output of another based on task-specific criteria such as relevance, depth, or creativity. Previous studies have explored both reference-based and reference-free variants of the LLM-as-a-judge paradigm for evaluating factual correctness. Reference-based frameworks such as DAFE (Badshah & Sajjad, 2025) and PoLL (Verga et al., 2024) show near-perfect alignment with human annotators when LLM judges are provided with gold-standard answers. However, this alignment deteriorates substantially when reference answers are excluded, highlighting a critical limitation: LLM-as-a-judge becomes unreliable without curated supervision (Ye et al., 2024; Kim et al., 2024; Huang et al., 2024). While reference-free approaches attempt to overcome this constraint, they inevitably rely on the model’s pre-trained knowledge and inherit its weaknesses—hallucination and bias. These findings underscore a fundamental challenge: how to enable factual, reference-free evaluation that grounds judgments in external, verifiable evidence rather than solely in a model’s internal knowledge.
In this paper, we present Tool-Augmented LLM Evaluation (TALE), a reference-free framework that bridges this gap by equipping LLM judges with the ability to actively gather and synthesize external evidence. Unlike conventional LLM-as-a-judge approaches that operate as closed systems, TALE implements an agent-as-a-judge paradigm (Zhuge et al., 2024) that iteratively generates web queries, retrieves information, reflects on findings, and refines its search strategy to verify or refute candidate responses. This active evidence-gathering process provides three key advantages: (1) it reduces dependency on the judge’s parameter knowledge, (2) it grounds evaluations in current, verifiable information, and (3) it enables assessment of novel or rapidly evolving topics where parameter knowledge might be outdated or incomplete.
Our key contributions include:
  • We introduce TALE, a novel framework that augments LLM-based evaluation with iterative, tool-assisted evidence gathering to overcome the limitations of both reference-based metrics and ungrounded LLM self-judgment.
  • We implement and evaluate a multi-step reasoning process, enabling LLM judges to actively verify factual claims rather than relying solely on internal knowledge
  • We demonstrate through extensive experiments that TALE is aligned with reference-based metrics such as F1 and achieves substantial to perfect agreements with human evaluators.
  • We provide ablation studies that quantify the impact of iterative refinement and query generation capabilities on evaluation performance. The results offer insights into the key components that drive TALE’s effectiveness.
2 Related Work
Evaluating LLMs is a critical yet challenging aspect of modern NLP research. We review existing approaches across the following categories:
2.1 Human and Reference-Based Evaluation
Human Evaluation.
Human assessment remains the gold standard for evaluating LLM-generated content (Chiang & Lee, 2023). However, it faces significant limitations: high cost (Zhou et al., 2023), limited reproducibility, subjective bias (Clark et al., 2021), and poor scalability for large-scale or continuous evaluation. These challenges have motivated the development of automated alternatives.
Reference-Based Metrics.
Traditional automatic evaluation relies on comparing model outputs against expert-annotated reference answers using metrics such as EM, F1, BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004). While efficient and reproducible, these approaches suffer from fundamental limitations: they cannot capture the diversity of valid responses, require costly reference annotations, and fail to adapt to evolving factual information. Recent work has attempted to address these limitations through learned metrics like BERTScore (Zhang et al., 2025) and BLEURT (Sellam et al., 2020). However, these metrics depend on the quality of reference answers and tend to overestimate the performance (Badshah & Sajjad, 2025).
2.2 LLM-as-a-Judge Approaches
The emergence of instruction-following capabilities in recent LLMs has enabled a new evaluation paradigm where models themselves serve as judges (Zheng et al., 2023). These approaches can be categorized based on their dependency on reference answers.
Reference-Based LLM Judges.
Several frameworks augment LLM judges with reference answers to improve evaluation reliability. DAFE (Badshah & Sajjad, 2025) provides LLM evaluators with gold-standard answers, achieving near-perfect alignment with human judgments in factual assessment. Similarly, PoLL (Verga et al., 2024), which employs a majority voting mechanism, demonstrates that instruction-tuned LLMs can effectively evaluate outputs when given reference answers. Auto-J (Li et al., 2023) combines reference answers with step-by-step reasoning to improve judgment accuracy. While these methods show promise, they inherit the fundamental limitations of requiring pre-defined references.
Reference-Free LLM Judges.
Addressing scalability concerns, several approaches eliminate reference dependency. G-Eval (Liu et al., 2023) implements direct evaluation by prompting models to assess outputs based on predefined criteria. Other methods include pairwise comparisons (Zheng et al., 2023), debate-style frameworks (Khan et al., 2024), and ensemble approaches (Zhang et al., 2024). These approaches have demonstrated success in subjective evaluation tasks like summarization or dialogue generation, where human preferences rather than factual correctness are the primary concern. However, when applied to factual evaluation, reference-free LLM judges often struggle with reliability (Ye et al., 2024; Kim et al., 2024), as they depend entirely on the judge model’s pre-trained knowledge and inherit its limitations in factual accuracy.
2.3 Tool-Augmented Evaluation
An emerging category attempts to overcome the limitations of closed-world LLM evaluators by incorporating external tools. WebGPT (Nakano et al., 2021) demonstrated how web-browsing capabilities improve factual correctness, while SelfCheckGPT (Manakul et al., 2023) used multiple independent generations to perform self-consistency checks. More closely related to our work, FActScore (Min et al., 2023) decomposes generated text into atomic facts and verifies each against retrieved external knowledge.
While these approaches leverage external information, they differ fundamentally from TALE in several aspects. First, most focus on evaluation through fact extraction and verification rather than holistic response assessment. Second, they typically perform single-pass evidence retrieval without the iterative refinement that characterizes TALE. Finally, they often lack explicit reasoning about what information to retrieve, instead relying on keyword-based or embedding similarity retrievals.
Our TALE framework addresses these gaps by implementing an Agent-as-a-judge paradigm (Zhuge et al., 2024) with iterative, adaptive evidence gathering, explicit reflection, and comprehensive reasoning, enabling more robust evaluation in scenarios where static reference answers are unavailable or insufficient.
3 Methodology
We introduce Tool-Augmented LLM Evaluation (TALE), a reference-free framework for evaluating LLM responses. Unlike conventional approaches that rely on fixed reference answers or human-annotated ground truths, TALE autonomously gathers and integrates external evidence to assess the correctness of free-form responses. Figure 1 illustrates the overall process of TALE.
Figure 1: Given the question, “What is the tallest building in the world?” and the candidate answer, “The tallest building in the world is the Burj Khalifa in Dubai.”, the process begins with an initial query generation based solely on the question to extract relevant information. This query is used for evidence gathering through web searches across multiple sources. Once the evidence is collected, it is processed by the summary module, which condenses it and extracts key insights. The summarized evidence is then passed to the reflector module, which assesses the sufficiency and relevance of the evidence. In case of insufficient or conflicting evidence, it triggers a query refinement step to improve the evidence through further iterations. After N iterations, the system enters the judgment phase where the judge synthesizes all available evidence and provides a final decision with rationale.
3.1 Candidate Response Generation
Let $x\in\mathcal{X}$ denote an input question (e.g., “What is the tallest building in the world?”). A candidate LLM $C$, such as GPT-3.5-turbo, produces a response $\hat{y}\in\mathcal{Y}$ according to $\hat{y}=C(x)$. This response $\hat{y}$ (e.g., “The tallest building in the world is the Burj Khalifa in Dubai”) is subsequently subject to evaluation via our tool-augmented process.
3.2 LLM-as-a-Judge
In our framework, an LLM is employed as a judge, denoted by $J$, to assess the correctness of $\hat{y}$ based on external evidence. Instead of relying on a fixed reference, the judge acts as an autonomous agent that leverages the aggregated evidence $E$ and reflection $R$, obtained through an iterative process, along with the input question $x$ and the candidate answer $\hat{y}$.
Formally, the judge produces a binary verdict $v\in\{0,1\}$ (i.e., True or False) and a brief natural language rationale $r_{J}$, such that:
$(v,r_{J})=J(x,\hat{y},E,R).$
Here, $J$ is a mapping defined as
$J:\mathcal{X}\times\mathcal{Y}\times\mathcal{E}\rightarrow\{0,1\}\times\mathbb{R},$
where $v=1$ indicates that the candidate answer $\hat{y}$ is deemed correct based on the evidence $E$, and $v=0$ otherwise. The judge gives precedence to evidence gathered from external sources but may fall back on their pre-trained knowledge when the evidence is incomplete or inconclusive.
3.3 Tool-Augmented Evaluation Process
TALE comprises the following modules:
Query Generation ($Q$):
In the first iteration ($i=1$), the query $q_{1}$ is constructed solely from the input question $x$: $q_{1}=Q(x)$. The goal is to retrieve background information relevant to the question domain, without assuming the correctness of any particular answer. For example, given the question “What is the capital of Switzerland?”, the model may generate a search query like “Capital of Switzerland” or “Switzerland official capital city”. This ensures that the evaluation process begins with a neutral, answer-agnostic evidence-gathering step.
In subsequent iterations ($i>1$), queries are refined using accumulated evidence and reflections stored in the short-term memory: $q_{i}=Q^{\prime}(\mathcal{M})$. The purpose is to resolve remaining uncertainty or retrieve more targeted information needed to verify or refute the candidate answer $\hat{y}$. For instance, if the candidate’s answer claims “Bern is the capital of Switzerland” but initial evidence suggests there might be a distinction between political and economic capitals, a refined query might be “Is Bern the official capital of Switzerland or is there another capital”.
Web Search ($S$):
The query $q_{i}$ is submitted to a web search via the Serper API, [^1] https://serper.dev/ to return real-time results. For each query, we retrieve the top $k=3$ search snippets or URLs, which serve as raw external evidence for the subsequent steps. The search results typically include titles, snippets, and source URLs, providing diverse perspectives from multiple information sources. This module enables TALE to access up-to-date and diverse information sources beyond the model’s pre-trained knowledge.
Summarization ($\Sigma$):
The retrieved search results—comprising multiple web snippets or document passages—are passed to a summarization module that condenses them into a focused evidence segment: $E_{i}=\Sigma(S(q_{i}))$. This module extracts salient factual content while filtering out redundant or irrelevant information. The resulting summary $E_{i}$ serves as a concise, interpretable knowledge unit that informs reflection. For example, a summary might state: “Multiple sources confirm that Bern is the de facto capital and seat of the federal government of Switzerland, though Switzerland does not have an official capital city designated by its constitution.” In TALE, this step is critical for managing context length and ensuring that only high-signal evidence is retained throughout the iterative reasoning process.
Reflection ($R$):
The summarized evidence $E_{i}$ is evaluated in relation to the input question $x$ and the candidate answer $\hat{y}$ to assess its relevance, sufficiency, and factual alignment. This step yields a reflection signal: $R_{i}=R(x,\hat{y},E_{i})$, which captures whether the current evidence supports, contradicts, or is inconclusive with respect to $\hat{y}$.
The reflection module also identifies missing information or ambiguities that can guide subsequent query refinement. By explicitly reasoning over the current evidence, reflection enables TALE to iteratively improve its evaluation behavior over multiple steps.
Short-Term Memory:
After each iteration, the tuple $(q_{i},E_{i},R_{i})$—consisting of the generated query, corresponding evidence, and reflection—is appended to the short-term memory buffer:
$\mathcal{M}\leftarrow\mathcal{M}\cup\{(q_{i},E_{i},R_{i})\}$
This memory $\mathcal{M}$ serves as an evolving trace of the evaluation process, accumulating context that informs subsequent query refinement and reasoning. The memory is append-only and scoped to a single evaluation episode, ensuring that each step builds upon the full history of prior actions, observations, and reflections.
Judge ($J$):
After $N$ iterations, the judge module synthesizes all gathered evidence to assess the correctness of the candidate answer. Specifically, it includes the aggregated evidence summary and reflections. Based on such prior components, $J$ produces a binary verdict $v\in\{0,1\}$ indicating whether $\hat{y}$ is factually correct, along with a natural language rationale $r_{J}$:
$(v,r_{J})=J(x,\hat{y},E_{\text{total}},R_{\text{total}}).$
The overall procedure is summarized in Algorithm 1.
Algorithm 1 TALE: Tool-Augmented LLM Evaluation
0:  Question $x$, candidate answer $\hat{y}$, number of iterations $N$
0:  Verdict $v\in\{0,1\}$ and rationale $r_{J}$
1:  Initialize short-term memory $\mathcal{M}\leftarrow[\,]$Memory buffer for query-evidence-reflection trace}
2:  $q_{1}\leftarrow Q(x)$Generate initial query from input question}
3:  $E_{1}\leftarrow\Sigma\bigl{(}S(q_{1})\bigr{)}$Retrieve and summarize external evidence}
4:  $R_{1}\leftarrow R(x,\hat{y},E_{1})$Reflect on evidence relevance and sufficiency}
5:  Append $(q_{1},E_{1},R_{1})$ to memory $\mathcal{M}$
6:  for $i\leftarrow 2$ to $N$ do
7:     $q_{i}\leftarrow Q^{\prime}\bigl{(}\mathcal{M}\bigr{)}$Refine query using accumulated memory}
8:     $E_{i}\leftarrow\Sigma\bigl{(}S(q_{i})\bigr{)}$
9:     $R_{i}\leftarrow R(x,\hat{y},E_{i})$
10:     Append $(q_{i},E_{i},R_{i})$ to memory $\mathcal{M}$
11:  end for
12:  $E_{\text{total}}\leftarrow\bigoplus_{i=1}^{N}E_{i}$Aggregate evidence summaries}
13:  $R_{\text{total}}\leftarrow\bigoplus_{i=1}^{N}R_{i}$Aggregate reflections}
14:  $(v,r_{J})\leftarrow J(x,\hat{y},E_{\text{total}},R_{\text{total}})$Final verdict and rationale}
15:  return  $(v,r_{J})$
4 Experimental Setup
This section details our experimental setup for evaluating TALE across multiple dimensions.
4.1 Models
In our experiments, we utilize: Gemini-1.5-pro (Team, 2024), GPT-3.5-turbo (Brown et al., 2020), and GPT-4o-mini (Team, 2023), both as candidates ($C$) and as judges ($J$) within the TALE framework, allowing us to assess both their ability to generate accurate responses and to evaluate responses from other models. Additionally, we also explore the potential of open-source LLMs as judges, including Mistral 7B [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). All experiments are conducted with a temperature of 00 to maximize determinism and reliability, as prior work demonstrates that higher temperatures degrade the performance of LLM-based evaluators (Hada et al., 2024). For brevity, we refer to these models as Mistral, Llama, Gemini, GPT-3.5, and GPT-4o throughout our analysis.
4.2 Datasets
We evaluate TALE on widely used free-form question-answering datasets that represent different question types, knowledge domains, and complexity levels. These includes AmbigQA (Min et al., 2020), HotpotQA (Yang et al., 2018), TriviaQA (Joshi et al., 2017), and Natural Questions (NQ-Open) (Kwiatkowski et al., 2019). Free-form question-answering underpins a broad range of practical applications, where maintaining accuracy and ensuring truthfulness are paramount (Gou et al., 2024). We also used FreshQA (Vu et al., 2023) to see TALE’s ability in detecting outdated knowledge. Due to computational constraints, we randomly sampled 300 instances from each dataset to ensure balanced representation across question types and difficulty levels. Each dataset provides reference answers that serve as ground truth for our reference-based baseline metrics (details in Appendix A.1).
4.3 Prompts
Our prompting strategy uses templates for both response generation and evaluation. For candidate models, we use few-shot Chain-of-Thought (CoT) prompts with 6 examples per dataset to elicit detailed, reasoning-based responses.
For the TALE framework, we design module-specific prompts that combine role instructions with step-by-step reasoning guidance. Each module (query generation, evidence summarization, reflection, refinement, and judgment) uses a one-shot prompt template that demonstrates the expected input-output behavior. These prompts are crafted to encourage explicit reasoning and evidence-focused analysis without assuming the correctness of any particular answer. Complete prompt templates and examples are provided in Appendix A.2.
4.4 Baselines
We compare TALE against several established evaluation approaches, including reference-based metrics and a reference-free judge. For the reference-based metrics, we implement two widely-used automatic metrics, Exact Match (EM) and F1. Additionally, we adopt a judge without tool-augmentation (G-Eval) as a baseline, following the approach from Liu et al. (2023), where the judge LLM evaluates candidate answers based solely on the question-answer pair without access to external tools or reference answers. In this setting, the judge relies entirely on its pre-trained knowledge to determine factual correctness. By maintaining the same judge model while removing the evidence retrieval mechanism, this baseline isolates the impact of tool augmentation (details in Appendix A.3).
Human Evaluation.
In addition to the above baselines, we recruited three graduate researchers with expertise in natural language processing to evaluate model outputs on AmbigQA and HotpotQA. Annotators were presented with input questions, corresponding reference answers, and anonymized model responses in a randomized order to prevent position or model identity bias. Each response was evaluated using a binary scoring system: 1 (“True”) for responses that accurately aligned with reference answers and demonstrated contextual relevance, and 0 (“False”) for responses that deviated from these criteria. The majority vote determined the final judgment. Appendix A.3 provides further details including annotation guidelines, percent agreement, and Fleiss’ Kappa scores.
4.5 Evaluation Metrics
To assess TALE’s performance, we used accuracy as the proportion of instances where the judge’s binary verdict (correct/incorrect) aligns with the ground truth derived from reference answers. For AmbigQA and HotpotQA that include human annotations, we calculate Cohen’s Kappa ($\kappa$) and Macro-F1 scores to evaluate TALE’s alignment with human majority votes, with Kappa accounting for agreement beyond chance and Macro-F1 addressing class balance. Additionally, we conduct ablation studies to quantify the impact of removing or modifying specific TALE components, using changes in agreement with human judgments as the primary measure. For component ablation studies, we measure the change in agreement with human judgments when removing or modifying specific TALE modules.
5 Results
This section briefly presents the results obtained from our experiments. We included additional results and analysis in the Appendix B.
By integrating external evidence into pre-trained knowledge, TALE significantly outperforms judges without tools.
Table 1 presents the raw performance of candidate LLMs using baseline methods and TALE. While judges without tools tend to overestimate accuracy by relying solely on pre-trained knowledge, TALE achieves far greater alignment with reference-based metrics like F1. For instance, GPT-3.5 as a judge within TALE reports an accuracy of 0.640 when evaluating itself on AmbigQA. This closely matches the F1 score (0.634) on the same task and model. In contrast, the same model without tool support drastically inflates its accuracy to 0.810. On the other hand, EM often underestimates performance by failing to account for valid paraphrases or alternative formulations. As a result, direct comparison between TALE and EM is often misleading.
TALE strongly correlates with human annotations
To evaluate alignment with human judgment, we compared TALE and baseline methods using majority vote annotations from three expert annotators on AmbigQA and HotpotQA. As depicted in Table 2, judges without access to reference answers or external tools rely heavily on their parameter knowledge and often confirm the candidate’s answer as correct. As a result, their agreement with human annotations is low, with Cohen’s kappa scores often below 0.40. In contrast, our proposed TALE framework achieves substantially higher agreement with human annotations. For instance, when GPT-4o evaluates itself without tools yields a $\kappa$ of only 0.375 on HotpotQA, while the same judge model under TALE reaches 0.701. Cohen’s $\kappa$ measures agreement beyond chance but can mislead under class imbalance, known as the kappa paradox. Therefore, we report Macro F1, which treats both positive and negative classes equally and provides a more balanced view of evaluation performance. As shown in Table 3, LLM-as-a-judge without access to tools or reference answers shows competitive macro F1 scores, but closer inspection reveals a tendency to over-predict correctness, leading to inflated recall at the expense of precision (see Section 5.1). In contrast, TALE consistently achieves the highest Macro F1 across models and tasks.
Candidate Task EM F1 Judge w/o Tool (Acc.) TALE (Acc.)
GPT-3.5 GPT-4o Gemini GPT-3.5 GPT-4o Gemini
GPT-3.5 AmbigQA 0.497 0.634 0.810 0.753 0.737 0.640 0.703 0.647
HotpotQA 0.337 0.474 0.857 0.760 0.700 0.503 0.543 0.527
NQ-Open 0.363 0.529 0.907 0.827 0.777 0.693 0.697 0.617
TriviaQA 0.743 0.810 0.887 0.863 0.817 0.813 0.850 0.777
GPT-4o AmbigQA 0.470 0.605 0.877 0.790 0.763 0.630 0.590 0.627
HotpotQA 0.343 0.465 0.863 0.770 0.860 0.500 0.480 0.530
NQ-Open 0.323 0.477 0.923 0.870 0.800 0.690 0.670 0.623
TriviaQA 0.763 0.837 0.930 0.900 0.857 0.853 0.867 0.800
Gemini AmbigQA 0.533 0.664 0.857 0.797 0.850 0.627 0.643 0.673
HotpotQA 0.347 0.495 0.833 0.787 0.753 0.503 0.507 0.547
NQ-Open 0.363 0.526 0.913 0.857 0.913 0.730 0.710 0.720
TriviaQA 0.793 0.861 0.910 0.917 0.893 0.870 0.877 0.817
Table 1: Raw performance of candidate LLMs obtained through different evaluators. Evaluation approaches include: (1) Reference-based metrics (EM, F1), (2) Judge w/o Tool (accuracy), and (3) TALE (accuracy).
TALE works better with more powerful LLMs.
TALE’s performance improves significantly when using more capable judge models. As shown in Tables 2 and 3, GPT-4o consistently outperforms GPT-3.5 and Gemini across both AmbigQA and HotpotQA. For instance, GPT-4o within TALE achieves a Cohen’s $\kappa$ of 0.701 on HotpotQA when evaluating GPT-4o as a candidate, compared to 0.540 with GPT-3.5 and 0.679 with Gemini. Similarly, GPT-4o reaches a macro F1 of 0.898 on AmbigQA when judging GPT-3.5, higher than GPT-3.5’s 0.784 and Gemini’s 0.830.
TALE can detect untruthful facts and outdated knowledge.
TALE excels at identifying false facts by cross-referencing claims with reliable external evidence. For example, when evaluating the question “Who sings the theme song for the show Half & Half?”, a candidate model incorrectly answered “Erica Campbell.” LLM judges without tools falsely evaluated the answer as correct. However, TALE retrieved accurate evidence confirming that “Melonie Daniels” performed the theme song. Using this information, TALE correctly rejected the candidate’s response, providing a clear rationale grounded in the evidence.
To evaluate TALE’s ability to detect outdated knowledge, we included FreshQA (Vu et al., 2023), Based on a random sample of 300 questions, we found that TALE consistently identified outdated information in candidate model responses. For instance, when asked “Where is EMNLP this year?”, candidate models often provided outdated responses based on their training data. TALE retrieved current information indicating the correct location, “Suzhou, China.” (see Appendix B.3).
TALE fixes incorrect reasoning traces.
We analyzed cases where candidate models produced logically inconsistent or unsupported reasoning. While some candidate responses contain plausible-sounding conclusions, their reasoning steps are often flawed or lack factual support. TALE’s reflection process enables it to detect these inconsistencies (see Table 8).
Candidate Task EM F1 Judge w/o Tool TALE
GPT-3.5 GPT-4o Gemini GPT-3.5 GPT-4o Gemini
GPT-3.5 AmbigQA 0.541 0.660 0.230 0.394 0.584 0.572 0.796 0.662
HotpotQA 0.596 0.755 0.156 0.258 0.348 0.413 0.557 0.525
GPT-4o AmbigQA 0.482 0.55 0.238 0.381 0.566 0.603 0.914 0.914
HotpotQA 0.544 0.662 0.169 0.375 0.363 0.540 0.701 0.679
Gemini AmbigQA 0.555 0.567 0.195 0.346 0.255 0.639 0.746 0.670
HotpotQA 0.490 0.662 0.172 0.272 0.265 0.633 0.693 0.594
Table 2: Cohen’s Kappa between the human majority votes and various evaluators. F1 scores are converted to binary using a $\tau=0.5$.
Candidate Task Judge w/o Tool TALE
EM F1 GPT-3.5 GPT-4o Gemini GPT-3.5 GPT-4o Gemini
GPT-3.5 AmbigQA 0.758 0.828 0.611 0.697 0.792 0.784 0.898 0.830
HotpotQA 0.790 0.877 0.527 0.609 0.665 0.706 0.779 0.762
GPT-4o AmbigQA 0.733 0.774 0.591 0.683 0.780 0.802 0.957 0.957
HotpotQA 0.763 0.830 0.532 0.669 0.675 0.770 0.850 0.839
Gemini AmbigQA 0.772 0.783 0.582 0.668 0.614 0.819 0.873 0.835
HotpotQA 0.732 0.829 0.553 0.617 0.619 0.816 0.846 0.797
Table 3: Macro F1 between the human majority vote and evaluators. F1 scores are converted to binary using a $\tau=0.5$.
5.1 Error analysis
To better understand the limitations of TALE in evaluating candidate responses, we conducted a manual error analysis. We randomly sampled 100 evaluation cases from the AmbigQA and HotpotQA datasets, focusing on instances where TALE disagreed with human annotators. We categorized the errors into the following categories: 1) Contextual misunderstanding: TALE generates inaccurate or incomplete queries when it misinterprets the intent of the candidate’s question. This is particularly evident in AmbigQA, where questions are often intentionally ambiguous or lack sufficient context, leading to the retrieval of irrelevant or contradictory evidence. 2) Incomplete evidence: TALE fails when the retrieved evidence is insufficient or lacks relevant information, specifically for recent events with limited online coverage. 3) Reasoning error: Despite accurate evidence, the judge model misinterprets the information or applies flawed reasoning. 4) Hallucination: In cases where evidence is ambiguous or inconclusive, TALE relies on its pre-trained knowledge, resulting in hallucinated rationales. 5) Conflicting evidence: In some cases, TALE encounters conflicting evidence across multiple search iterations. While the framework is designed to iteratively refine its understanding and gather reliable information, judges sometimes over-rely on earlier sources or fail to appropriately weigh the credibility of conflicting information (details in Appendix B.5).
5.2 Effect of iterations
We conducted an ablation study to evaluate the effect of the number of iterations in the TALE framework. Figure 2 shows that increasing the number of iterations generally improves the judge performance across candidate models on both AmbigQA and HotpotQA. Iteration 3, which serves as our default configuration, consistently provides the best trade-off between performance and computational cost. However, performance declines at iteration 4 across all models. This drop is likely caused by an overabundance of sources, resulting in the accumulation of redundant or irrelevant information, increased context length, and potential model confusion. These factors can dilute decision quality and lead to marginal or negative returns.
5.3 Effect of removing the query making
To evaluate the importance of the query generation module in TALE, we conduct an ablation study by removing the query-making step and directly using the input question for evidence retrieval. Table 4 shows that TALE consistently outperforms the query-free baseline. The most notable improvements are observed on AmbigQA, where TALE achieves a Macro F1 of 0.9568 compared to 0.8438 without query making. In HotpotQA, while the performance gain is smaller, TALE still demonstrates clear advantages, particularly due to its ability to adaptively generate focused queries that facilitate multi-hop reasoning.
Figure 2: Left: The effect of iterations for AmbigQA. Right: The effect of iterations for HotpotQA. GPT-4o is used as a judge here.
Candidate Task Judge w/o Query (GPT-4o) TALE (GPT-4o)
GPT-3.5 AmbigQA 0.8402 0.8978
HotpotQA 0.7632 0.7786
GPT-4o AmbigQA 0.8438 0.9568
HotpotQA 0.8233 0.8500
Gemini AmbigQA 0.8466 0.8725
HotpotQA 0.7993 0.8457
Table 4: Comparison of Macro F1 scores between Judge w/o Query (GPT-4o as-as-judge) and TALE (GPT-4o as-as-judge). Note that directly using the input question without generating refined queries is still considered a form of querying, but for clarity, we refer to this setting as w/o Query.
6 Conclusion
We presented TALE, a novel framework for evaluating LLMs. TALE integrates external evidence via an iterative process. Through extensive experiments, we found that TALE achieves substantial to perfect agreement with human annotators. TALE is also aligned with reference-based metrics such as F1. However, studying TALE compared to human evaluations, we found that TALE is a reliable alternative to reference-based metrics with interpretable evidence-aware evaluations. Regardless, TALE has limitations. Its short-term memory provides a persistent record of the evaluation process. While this enhances interpretability and traceability, it remains constrained by the model’s ability to process information within its context window during subsequent reasoning. Future work could explore integrating a long-term memory with a recall-based mechanism that selectively retrieves relevant traces from past evaluations, such as episodic and semantic memory (Park et al., 2023). In addition to this, TALE may inherit biases from external data sources, including sycophancy bias (Sharma et al., 2023), where the model may align with false claims instead of critically assessing them. Addressing these biases and ensuring robust evidence selection will be essential for broader deployment.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI), and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada.
Appendix A Experimental detail
A.1 Datasets
We evaluate TALE on widely used free-form question-answering datasets that represent different question types, knowledge domains, and complexity levels. Free-form question-answering underpins a broad range of practical applications, where maintaining accuracy and ensuring truthfulness are paramount (Gou et al., 2024). Evaluating large-scale datasets can be computationally prohibitive; therefore, we randomly sample 300 instances from each dataset, ensuring a balanced representation across question types and difficulty levels. This sampling strategy provides a fair evaluation while maintaining computational feasibility. Each dataset provides reference answers that serve as ground truth for our reference-based baseline metrics, allowing us to compare the performance of TALE against established evaluation approaches. Our selected datasets are:
AmbigQA
(Min et al., 2020) Contains questions with multiple valid answers due to inherent ambiguities, challenging evaluators to consider multiple interpretations.
HotpotQA
(Yang et al., 2018) Features multi-hop reasoning questions that require synthesizing information from multiple sources.
Natural Questions (NQ-Open)
(Kwiatkowski et al., 2019) Consists of real user queries from Google Search, representing naturally occurring information needs.
TriviaQA
(Joshi et al., 2017) Includes trivia questions from various domains, testing breadth of knowledge and factual recall.
FreshQA
(Vu et al., 2023) Contains questions about recent events occurring after most LLMs’ training cutoff, specifically designed to test knowledge updating capabilities.
A.2 Prompting
We employ a template-based prompting strategy for both response generation and evaluation. For candidate models, we utilize few-shot Chain-of-Thought (CoT) prompts, incorporating 6 examples per dataset to encourage detailed, reasoning-driven, and structured responses (see Figure 3).
Candidate Answer Generation Instructions: Answer the following questions. For each question, provide a complete answer, and conclude with the phrase: "So the answer is: [Answer]." Q: What is the elevation range for the area that the eastern sector of the Colorado orogeny extends into? A: Let’s think step by step. The eastern sector of Colorado orogeny extends into the High Plains. High Plains rise in elevation from around 1,800 to 7,000 ft. So the answer is: 1,800 to 7,000 ft. Q: Musician and satirist Allie Goertz wrote a song about the "The Simpsons" character Milhouse, who Matt Groening named after who? A: Let’s think step by step. Milhouse was named after U.S. president Richard Nixon. So the answer is: Richard Nixon. Q: Which documentary is about Finnish rock groups, *Adam Clayton Powell* or *The Saimaa Gesture*? A: Let’s think step by step. *Adam Clayton Powell* is a documentary about an African-American politician, not Finnish rock groups. So the documentary about Finnish rock groups must instead be *The Saimaa Gesture*. So the answer is: The Saimaa Gesture. Q: What profession do Nicholas Ray and Elia Kazan have in common? A: Let’s think step by step. The professions of Nicholas Ray are director, screenwriter, and actor. The professions of Elia Kazan are director, producer, screenwriter, and actor. So the professions Nicholas Ray and Elia Kazan have in common are director, screenwriter, and actor. So the answer is: director, screenwriter, actor. Q: Which magazine was started first, *Arthur’s Magazine* or *First for Women*? A: Let’s think step by step. *Arthur’s Magazine* was started in 1844. *First for Women* was started in 1989. 1844 (Arthur’s Magazine) < 1989 (First for Women), so *Arthur’s Magazine* was started first. So the answer is: Arthur’s Magazine. Q: Were Pavel Urysohn and Leonid Levin known for the same type of work? A: Let’s think step by step. Pavel Urysohn is a mathematician. Leonid Levin is a mathematician and computer scientist. So Pavel Urysohn and Leonid Levin have the same type of work. So the answer is: Yes.
Figure 3: Examples of few-shot Chain-of-Thought (CoT) prompts for candidate answer generation.
TALE prompts candidate LLMs using few-shot Chain-of-Thought (CoT) reasoning to elicit faithful and interpretable outputs across each module. Each module in the evaluation pipeline is guided by a carefully constructed prompt, often in a one-shot format, that combines role-playing instructions with explicit reasoning goals. Below, we describe the prompting strategy for each component.
Query Generation.
The query generation module converts an input question $x\in\mathcal{X}$ into an initial search query without referencing the candidate answer. The prompt instructs the model to reflect step-by-step on the most relevant aspects and keywords before proposing a final query.
Query generation Your goal is to generate a targeted web search query. Before you produce the final query, think carefully about: 1. The question’s key concepts or keywords (e.g., important names, dates). 2. Whether the question might be ambiguous or reference multiple possible answers (e.g., a book with the same title by different authors, or a modern text about a historical figure). Question: {question} Return your response as a JSON object with ALL three exact keys: "query": The search query string. "aspect": The specific aspect of the question to focus on. "rationale": A brief explanation of why this query is relevant, including your chain-of-thought reasoning. Example Output:
  "query": "Apollo 11 moon landing year + NASA + 1969",
  "aspect": "historical event",
  "rationale": "The question asks about Apollo 11’s landing year,
                so I’m including NASA, year, and 1969 to get relevant info."
Figure 4: Example of a query generation prompt and output used for the initial query generation step.
Evidence Summarization.
To reduce raw search results $S(q_{i})$, the summarization module uses a Chain-of-Thought (CoT) prompt that walks the model through evaluating and synthesizing relevant content. The prompt emphasizes factual grounding and asks the model to avoid repetition and speculation.
Evidence summarization You are a summarization assistant. Carefully review the raw search results and then provide a concise summary of the key information relevant to the question. Raw Search Results: {raw_results} Return your summary as plain text. Keep it neutral and focused on the question. If results conflict, mention that briefly. Do not add extra commentary. Example Output (Plain Text):
"Result 1 says X about the event date,
 Result 2 says Y but doesn’t mention the exact date.
 Overall, it references 1969."
Figure 5: Example prompt used in the evidence summarization step, guiding the model to generate a concise, unbiased summary from raw search results.
Iterative Reflection.
The reflection module analyzes the current evidence summary $E_{i}$ in relation to the input question $x$ and candidate answer $\hat{y}$. The prompt guides the model to assess whether the evidence supports, contradicts, or is inconclusive with respect to the answer. It also encourages identifying what information is missing.
Example prompt for iterative reflection You are a research assistant tasked with analyzing the gathered evidence in relation to the question and candidate answer. Think step by step—explain your reasoning and note any gaps or additional details that might be needed. Do not provide a final decision; simply offer your chain-of-thought reflection. Question: {question} Candidate Answer: {candidate_answer} Evidence Summary: {evidence_summary} Return your response as a JSON object with a single key: "reflection": Your chain-of-thought reflection summarizing your analysis. Example Output:
  "reflection": "I observed that the evidence overwhelmingly confirms
                 that Apollo 11 landed on the moon in 1969, though there
                 is slight variation in the reported landing times across
                 sources. Additional authoritative sources might help
                 resolve these minor discrepancies."
Figure 6: Example prompt used in the iterative reflection step, instructing the model to analyze the relationship between the question, candidate answer, and evidence summary.
Query Refinement.
To improve the evidence retrieved in future steps, the query refinement module generates a new query by analyzing the short-term memory contents—specifically, the previous query $q_{i}$, evidence $E_{i}$, and reflection $R_{i}$. The Chain-of-Thought (CoT) prompt instructs the model to identify remaining uncertainties or gaps and generate a refined, more targeted query.
Example prompt for query refinement You are a research assistant. Before refining the search query, analyze the existing evidence and reflect on what keywords might be missing or need emphasis. Think step by step and then produce your final refined query. Question: {question} Current Search Query: {current_query} Aggregated Evidence Summary: {evidence_summary} Iterative Reflection: {iterative_reflection} If the evidence still does not resolve the question or if there might be an alternative perspective, incorporate additional, more specific keywords to explore those possibilities. For instance: Add relevant dates or historical context. Use synonyms or alternate phrasings for ambiguous or repeated terms. Specify a domain or subject area (e.g., “film,” “novel,” “historical figure”) if it reduces confusion. Highlight the location, time period, or any unique aspect not yet included in the current query. Return your response as a JSON object with ALL three exact keys: "query": The refined search query. "aspect": The specific aspect being targeted with the refined query. "rationale": A brief explanation of your reasoning (chain-of-thought) and why this refinement is needed. Example Output:
  "query": "Apollo 11 detailed timeline moon landing 1969",
  "aspect": "chronological sequence",
  "rationale": "The initial query did not specify the temporal progression
                of events. I refined it to target a detailed timeline of
                the Apollo 11 mission in 1969 to capture the sequence of
                key events."
Figure 7: Example prompt for query refinement, guiding the model to analyze evidence and iteratively generate a more targeted query.
Judgment.
Finally, the judgment module evaluates whether the candidate’s answer $\hat{y}$ is factually correct based on the accumulated external evidence.
The Chain-of-Thought (CoT) prompt instructs the model to reason step-by-step using the evidence and produce a binary decision, True/False decision, along with rationale.
In cases where the evidence is insufficient or contradictory, the prompt explicitly instructs the model to either defer to its prior knowledge or explain uncertainty. The output is formatted as a JSON object with keys "decision" and "explanation." All modules are executed within a single LLM agent under a unified prompting interface.
Prompt to the judge model You are a critical evaluator. You have:
1. The question and the candidate answer,
2. The evidence summary from multiple iterative searches (these may sometimes contain overlapping or conflicting info), 3. The chain-of-thought reflection from prior steps, 4. Your own broad knowledge (only if the above are inconclusive). Follow these guidelines:
- If the summarized evidence and reflections strongly conflict with the candidate answer, conclude "False."
- If the evidence strongly confirms the candidate answer, conclude "True." - If the evidence is inconclusive or incomplete, but your own knowledge supports the answer, you may conclude "True" if confident. Otherwise, conclude "False" or state insufficient information. - When the retrieved evidence is irrelevant, prioritize the chain-of-thought reflections and your own knowledge. Produce your conclusion in JSON with: - "decision": "True" or "False" - "explanation": A concise reason (including your step-by-step reasoning) describing how you arrived at the verdict. Input: Question: {question} Candidate Answer: {candidate_answer} Evidence Summary: {evidence_summary} Reflection: {reflection} Example Output:
  "decision": "True",
  "explanation": "The evidence overwhelmingly confirms that Apollo 11
                  landed on the moon in 1969. While minor discrepancies
                  exist in the reported times, they do not undermine the
                  main conclusion. Additional verification is unnecessary."
Figure 8: Example prompt for the judgment step, instructing the model to analyze evidence and reflections to generate a final verdict with justification.
A.3 Baselines
We compare TALE against several established evaluation approaches:
A.3.1 Reference-Based Metrics.
We implement two widely-used automatic metrics that rely on comparison with dataset-specific reference answers:
  • Exact Match (EM) measures whether the model’s answer exactly matches any of the reference answers after normalization.
  • F1 Score computes the harmonic mean of precision and recall between the token sets of the model’s answer and the references, providing a softer measure of overlap.
A.3.2 Judge without Tool-Augmentation
Following the approach from Liu et al. (2023), we implement a reference-free baseline where the judge LLM evaluates candidate answers based solely on the question-answer pair, without access to external tools or reference answers. The judge relies entirely on its pre-trained knowledge to determine factual correctness. This baseline isolates the impact of tool augmentation in TALE by maintaining the same judge model while removing the evidence retrieval mechanism.
A.3.3 Human Evaluation
We recruited three graduate researchers with expertise in natural language processing to evaluate model outputs on AmbigQA and HotpotQA. Annotators were presented with input questions, corresponding reference answers, and anonymized model responses in a randomized order to prevent position or model identity bias. Each response was evaluated using a binary scoring system: 1 (“True”) for responses that accurately aligned with the reference answers and demonstrated contextual relevance, and 0 (“False”) for responses that deviated from these criteria.
Evaluation Rationale
Due to budget and resource constraints, we focused our human evaluation on AmbigQA and HotpotQA. These datasets were chosen because they represent challenging real-world scenarios involving multi-hop reasoning and ambiguous question-answering, making them ideal for assessing TALE’s effectiveness. Evaluating additional datasets would have significantly increased the time and cost of human annotations.
Furthermore, we evaluated 300 randomly sampled instances from each dataset, resulting in 600 samples per model across the two tasks. With three candidate models, the total number of samples evaluated was resulted in 1,800. Conducting large-scale human evaluation beyond this would incur substantial annotation costs and additional cognitive load on annotators. By limiting the sample size, we maintained a balance between evaluation comprehensiveness and resource efficiency.
Evaluation Guidelines
To ensure consistent assessments, annotators followed the guidelines inspired by established evaluation protocols. Annotators were instructed to evaluate responses based on the following principles:
  • Semantic equivalence: A response is marked True if it conveys the same core information as the reference answer, even if phrased differently using synonyms, paraphrasing, or structural variations. Additional contextual information is acceptable as long as it is factually correct and does not alter the original meaning.
  • Factual Accuracy: Responses that contain factual errors, omit essential information, or introduce misleading content are marked False. If a response partially answers the question but excludes critical elements, it is considered incorrect.
  • Multiple Reference Answers: In cases with multiple reference answers, a response is deemed correct if it is fully aligned with at least one reference.
  • Fact-Checking: Annotators are allowed to consult external resources, such as search engines or online encyclopedias, to verify specific facts when uncertain. However, the reference answers served as the primary benchmark for correctness.
  • Documenting Ambiguity: Annotators are encouraged to document cases where the evaluation is uncertain or requires further clarification. These cases were discussed collaboratively to ensure consensus.
By adhering to these guidelines, we ensured reliable and consistent human evaluations.
Inter Human Annotator Agreement
We calculated Fleiss’ Kappa ($\kappa$) and percent agreement to measure inter-annotator agreement.
Fleiss’ Kappa is defined as:
$\kappa=\frac{\bar{P}-P_{e}}{1-P_{e}},$
where $\bar{P}$ is the average observed agreement among annotators, and $P_{e}$ is the expected agreement by chance.
Percent agreement is calculated as:
Percent Agreement=Number of AgreementsTotal Number of $\text{Percent Agreement}=\frac{\text{Number of Agreements}}{\text{Total Numberof Annotations}}\times 100.$
A.4 Evaluation Metrics
To assess TALE’s performance, we use multiple evaluation metrics:
Accuracy.
We measure the proportion of instances where the judge’s binary verdict (correct/incorrect) aligns with the ground truth derived from reference answers.
Agreement with Human Judgment.
For the AmbigQA and HotpotQA subsets with human annotations, we calculate Cohen’s Kappa ($\kappa$), majority voting, and Macro-F1 scores to assess agreement between TALE’s verdicts and human majority votes. These metrics were chosen because they account for both agreement beyond chance ($\kappa$) and class balance (Macro-F1).
Cohen’s Kappa: Cohen’s Kappa measures the agreement between two annotators while correcting for chance agreement. It is defined as:
$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}},$
where $P_{o}$ is the observed agreement, and $P_{e}$ is the expected agreement by chance.
Majority Voting: In majority voting, the final decision is determined based on the majority of annotators’ labels. Given $n$ annotators and a binary classification, the majority label is defined as:
$y_{\text{majority}}=\begin{cases}1&\text{if }\sum_{i=1}^{n}y_{i}>\frac{n}{2},\\ 0&\text{otherwise},\end{cases}$
where $y_{i}$ represents the label assigned by the $i$th annotator.
Macro F1 Score: Macro F1 evaluates the balance between precision and recall for each class and averages the results. It is calculated as:
[…]
where $C$ is the number of classes, and $\text{Precision}_{c}$ and $\text{Recall}_{c}$ are the precision and recall for class $c$.
Appendix B Additional results
In this section, we included additional results obtained through our experiments.
B.1 Inter-human annotator agreement
Table 5 presents the human annotator agreement results for the AmbigQA and HotpotQA across three candidate models. The results indicate consistently high agreement among annotators.
Task Model Percent Agreement (%) Fleiss’ Kappa Samples
AmbigQA GPT-3.5 98.3 0.972 300
GPT-4 98.3 0.976 300
Gemini 97.0 0.953 300
HotpotQA GPT-3.5 98.3 0.978 300
GPT-4 98.3 0.978 300
Gemini 98.3 0.977 300
Table 5: Human annotator agreement results on AmbigQA and HotpotQA tasks.
B.2 TALE with a small open-source model
We used Mistral 7B to investigate how smaller open-source models perform within our TALE framework. Specifically, we employed Mistral 7B as a judge in TALE to evaluate candidate GPT-3.5 on the AmbigQA and HotpotQA datasets. This evaluation provides insights into the effectiveness of smaller models in assessing complex reasoning and factual correctness.
Table 6 illustrates that Mistral 7B, despite its smaller size, demonstrates a reasonable capability in evaluating complex reasoning and factual correctness. On AmbigQA, Mistral 7B achieves a Cohen’s Kappa of 0.5910 and a Macro F1 of 0.7955, indicating a moderate agreement. On HotpotQA, which involves multi-hop reasoning, its Cohen’s Kappa of 0.3282 and Macro F1 of 0.6617 suggest greater challenges in assessing factual accuracy and reasoning depth.
However, a notable limitation of Mistral 7B is its frequent difficulty in following instructions precisely, particularly when handling complex queries that require deep understanding. Additionally, its smaller context window limits its ability to maintain coherence across long reasoning chains. These issues are evident in scenarios where the model fails to properly parse iterative reflection responses or refines search queries incorrectly. Furthermore, Mistral 7B sometimes generates irrelevant reflections or fails to recognize when no supporting evidence is available, leading to errors in judgment. Despite these challenges, Mistral 7B remains a valuable option for resource-constrained environments where efficient evaluation is prioritized over peak accuracy (Badshah & Sajjad, 2024).
Candidate Model Task Accuracy Cohen’s Kappa Macro F1
GPT-3.5 AmbigQA 0.74 0.5910 0.7955
GPT-3.5 HotpotQA 0.62 0.3282 0.6617
Table 6: Mistral 7B as a judge within the TALE evaluating GPT-3.5 answers for AmbigQA and HotpotQA datasets.
B.3 TALE can detect untruthful facts and outdated knowledge.
TALE’s iterative evidence-gathering and reflection process enables it to detect untruthful claims and identify outdated information. By continuously refining its search queries and critically evaluating retrieved evidence, TALE can distinguish between correct and incorrect candidate answers, even when the misinformation is subtle. This capability is particularly valuable in dynamic domains where factual knowledge changes over time.
Table 7 presents an example where a candidate answer incorrectly claims that the last perfect game in Major League Baseball was thrown by Félix Hernández in 2012. Through iterative search and reflection, TALE discovers recent evidence confirming that Domingo Germán pitched a perfect game in 2023, successfully identifying the outdated information and concluding that the candidate’s answer is incorrect.
Component Content
Question When was the last perfect game thrown in baseball?
Candidate answer The last perfect game in Major League Baseball was thrown by Félix Hernández of the Seattle Mariners on August 15, 2012.
Initial query last perfect game in MLB history
Evidence summary The last perfect game in MLB history was thrown by Domingo Germán of the New York Yankees on June 28, 2023, against the Oakland Athletics.
Reflection The candidate answer states that the last perfect game was thrown by Félix Hernández in 2012. However, the evidence confirms that the last perfect game was actually thrown by Domingo Germán in 2023. This indicates the candidate answer is outdated.
Refined duery perfect game baseball MLB after 2012
Final decision False
Explanation The candidate answer states that Félix Hernández threw the last perfect game on August 15, 2012. However, all evidence confirms that Domingo Germán pitched a perfect game on June 28, 2023. This makes the candidate answer outdated and incorrect.
Table 7: An example where TALE detects an outdated fact in the candidate answer by identifying recent evidence through iterative search and reflection.
We further evaluated TALE’s performance using FreshQA (Vu et al., 2023). In this evaluation, we used GPT-4o as the judge model within TALE to assess the accuracy of candidate responses generated by GPT-3.5. The results demonstrate that TALE performed notably well in these contexts, achieving an agent-based raw accuracy of 38.33%, which is significantly closer to reference-based metrics compared to its EM score of 25.00% and F1 score of 35.40%. This highlights TALE’s strength in adapting to evolving information and accurately identifying untruthful or outdated claims. From these raw scores, it is also evident that a pre-trained model like GPT-3.5 often struggles to accurately respond to factual questions in rapidly evolving domains.
B.4 TALE fixes incorrect reasoning traces.
TALE’s iterative search and reflection process enables it to identify and correct flawed reasoning in candidate answers. Even when a final answer is correct, the candidate’s reasoning may contain factual errors. By refining its search queries and critically analyzing the evidence, TALE can highlight such errors and provide a more accurate rationale.
Table 8 presents an example where the candidate’s answer correctly concludes that Sherwood Stewart was born before Javier Frana. However, the reasoning contains a factual inaccuracy, falsely stating Stewart’s birth year as 1957 instead of the correct 1946. Through multiple iterations, TALE gathers evidence to correct this mistake while maintaining the correct conclusion.
Component Content
Question Who was born first, Javier Frana or Sherwood Stewart?
Candidate answer Let’s think step by step. Javier Frana was born in 1966. Sherwood Stewart was born in 1957. 1957 (Sherwood Stewart) < 1966 (Javier Frana), so Sherwood Stewart was born first. So the answer is: Sherwood Stewart.
Search query Javier Frana birthday
Evidence summary Javier Frana was born on December 25, 1966. No evidence was found regarding Sherwood Stewart’s birthdate.
Reflection The candidate answer claims that Sherwood Stewart was born in 1957, but the evidence only confirms Javier Frana’s birth year. Further search is required to verify Stewart’s actual birthdate.
Refined query (Iteration 2 Sherwood Stewart tennis birthdate
Evidence summary (iteration 2) Sherwood Stewart was born on June 6, 1946, according to multiple sources. Javier Frana was born on December 25, 1966.
Reflection (iteration 2) While the candidate answer initially provides an incorrect birth year for Stewart (1957), the evidence confirms that he was actually born in 1946. The candidate’s final conclusion that Stewart was born first remains correct despite the factual error.
Final decision True
Explanation The candidate answer correctly identifies Sherwood Stewart as being born first. Although it initially provided an incorrect birth year for Stewart, the evidence from iterations 2 and 3 confirms the correct birthdate as 1946. Therefore, the final answer is correct, but the reasoning was flawed.
Table 8: An example where TALE detects and corrects an incorrect reasoning trace. While the candidate’s answer concludes correctly, the system highlights the factual inaccuracy in the intermediate reasoning.
B.5 TALE failure cases
To better understand the limitations of TALE in evaluating candidate responses, we conducted a manual error analysis. We randomly sampled 100 evaluation cases from the AmbigQA and HotpotQA datasets, focusing on instances where TALE disagreed with human annotators. The errors were categorized into different categories, with representative examples presented in the Tables.
Contextual Misunderstanding.
TALE generates inaccurate or incomplete queries when it misinterprets the intent of the candidate’s question. This is particularly evident in AmbigQA, where questions are often intentionally ambiguous or lack sufficient context, leading to the retrieval of irrelevant or contradictory evidence.
Table 9 presents an example where TALE faced a contextual misunderstanding. The question asked about the opening date of the Royal Liverpool Hospital before redevelopment. However, the candidate answer referred to the hospital’s reopening in 2022, following redevelopment, rather than its original opening date in 1978. Despite iterative search and reflection, TALE failed to align with the actual intent of the question, leading to a false final decision.
Component Content
Question When does the new Royal Liverpool Hospital open before redevelopment?
Candidate answer The new Royal Liverpool Hospital is scheduled to open in 2022, following delays in construction and redevelopment. So the answer is: 2022.
Search query Royal Liverpool Hospital original opening date before redevelopment.
Evidence summary The new Royal Liverpool Hospital opened in 1978 after services were transferred from the old Liverpool Royal Infirmary.
Reflection The candidate answer claims that the new hospital was scheduled to open in 2022. However, the evidence confirms its original opening was in 1978. The judge misinterpreted the question’s intent, mistaking the redevelopment’s reopening for the original opening date.
Refined query (Iteration 2) Royal Liverpool Hospital redevelopment opening date 2022.
Evidence summary (Iteration 2) Structural repairs were completed, allowing the hospital to reopen in 2022. The move to the new hospital began on September 28, 2022.
Reflection (Iteration 2) The evidence confirms the hospital reopened in 2022, but the question specifically asks about the original opening before redevelopment. Despite the accurate evidence about the redevelopment timeline, the candidate answer misaligned with the question intent.
Final decision True (Incorrect)
Explanation TALE concluded that the candidate answer was correct, accepting the 2022 reopening as the answer. However, the correct answer was 1978, referring to the hospital’s original opening. This was a case of contextual misunderstanding.
Table 9: Example of a contextual misunderstanding where TALE misinterpreted the question’s intent by focusing on the hospital’s reopening in 2022 instead of its original opening in 1978.
Incomplete Evidence.
TALE fails when the retrieved evidence is insufficient or lacks relevant information, especially for verifying specific factual claims (see Table 10).
Component Content
Question Wing Chun starred Michelle Yeoh and a Hong Kong actor, martial artist, and film director who won what tournament multiple times?
Candidate answer The Hong Kong actor in Wing Chun is Donnie Yen. Donnie Yen won the multiple times wushu tournament champion. So the answer is: wushu tournament.
Evidence summary The film "Wing Chun," released in 1994, stars Michelle Yeoh and Donnie Yen, directed by Yuen Woo-ping. The search results do not provide evidence that he won multiple wushu tournaments.
Reflection The evidence summary lacks any mention of specific wushu tournament victories by Donnie Yen. The absence of evidence to validate the candidate’s claim indicates incomplete information. Additional reliable sources would be required to confirm the accuracy of the answer.
Final decision False
Table 10: Example of incomplete evidence where TALE could not retrieve sufficient information to verify the candidate’s claim.
Reasoning Error.
Despite accurate evidence, the judge model misinterprets the information or applies flawed reasoning. Reasoning errors often occur when the model fails to apply appropriate logic to the available evidence or when it misrepresents the intent of the question. Table 11 illustrates a case where TALE incorrectly concluded that Guglielmo Marconi was the first to achieve wireless telegraphy communication. The evidence suggested that while Marconi developed the first practical system, earlier demonstrations, including those by Jagadish Chandra Bose, may have preceded it.
Component Content
Question Who made communication by wireless telegraphy for the first time?
Candidate answer The first successful communication by wireless telegraphy was achieved by Guglielmo Marconi. He is credited with developing the first practical system for wireless telegraphy in the late 19th century. So the answer is: Guglielmo Marconi.
Search query first successful wireless telegraphy transmission + inventor + date
Evidence summary Guglielmo Marconi is known for creating a practical radio wave-based wireless telegraph system. He began working on wireless telegraphy in the 1890s.
Reflection The candidate answer asserts that Guglielmo Marconi made the first successful wireless telegraphy communication. While the evidence supports Marconi’s development of the first practical wireless telegraph system, the question explicitly asks who made the first communication, not the first practical one. Further evidence is required to investigate whether anyone transmitted information wirelessly before Marconi.
Refined query (Iteration 2) pre-Marconi wireless experiments OR early electromagnetic communication OR history of radio waves before 1900 OR wireless telegraphy before Marconi
Evidence summary (Iteration 2) While Marconi is credited with the first practical wireless telegraphy system, there were earlier demonstrations of electromagnetic wave transmission, including Jagadish Chandra Bose’s 1895 experiment. Wireless telegraphy was also reported on ships in 1891.
Reflection (Iteration 2) The evidence indicates the existence of earlier wireless communication attempts, particularly by Jagadish Chandra Bose. Despite this, the candidate answer incorrectly concluded that Marconi was the first without addressing these earlier cases. The system failed to recognize the distinction between "first communication" and "first practical system."
Final decision False (incorrect)
Explanation While the evidence confirms that Guglielmo Marconi developed the first practical wireless telegraphy system, the evidence also suggests prior communications using wireless methods. The distinction between a practical system and the first-ever communication was not correctly reasoned, leading to an incorrect final decision.
Table 11: Example of a reasoning error where TALE misinterpreted the question’s intent, failing to account for earlier instances of wireless telegraphy communication.
Hallucination.
In cases where evidence is unclear or inconclusive, TALE relies on its pre-trained knowledge, resulting in hallucinated rationales. Hallucination errors often occur when the model confidently asserts false information without sufficient evidence. Table 12 illustrates a case where TALE incorrectly concluded that Sarah Michelle Gellar, a cast member of *Scream 2*, was a star of the TV show *Roseanne*, despite the absence of supporting evidence.
Component Content
Question Which "Roseanne" star is in Scream 2?
Candidate answer Let’s think step by step. "Roseanne" star in Scream 2 is Sarah Michelle Gellar. So the answer is: Sarah Michelle Gellar.
Search query Roseanne Barr Scream 2 cast
Evidence summary Laurie Metcalf, who played Jackie Harris on the TV series Roseanne, also appeared in the film Scream 2 in 1997. Sarah Michelle Gellar, while a star in Scream 2, is not a Roseanne star.
Reflection The candidate answer incorrectly claims that Sarah Michelle Gellar was a star on Roseanne. This statement is a hallucination as no evidence supports this claim. The evidence correctly identifies Laurie Metcalf as the only Roseanne star in Scream 2.
Refined query (Iteration 2) Roseanne cast members in Scream 2
Evidence summary (Iteration 2) Laurie Metcalf, who starred as Jackie Harris in Roseanne, is confirmed to be in Scream 2. Sarah Michelle Gellar also appeared in the film, but she has no connection to Roseanne.
Reflection (Iteration 2) The evidence confirms Laurie Metcalf’s participation in both Roseanne and Scream 2. Despite further evidence confirming no connection between Gellar and Roseanne, the model continued to assert a false claim.
Final decision False (hallucination)
Explanation The candidate answer confidently asserted that Sarah Michelle Gellar was a Roseanne star, despite all evidence indicating that Laurie Metcalf was the only Roseanne star in Scream 2. This case is a clear example of hallucination, where the model generated a false fact with misplaced confidence.
Table 12: Example of hallucination where TALE falsely identified Sarah Michelle Gellar as a star of Roseanne.
Conflicting Evidence.
In some cases, TALE encounters conflicting evidence across multiple search iterations. While the framework is designed to iteratively refine its understanding and gather reliable information, it sometimes struggles to appropriately resolve contradictory information. Table 13 presents a scenario where the model failed to handle conflicting evidence when assessing the number of band members in Lush and P.O.D.
Component Content
Question Do Lush and P.O.D. both consist of four band members?
Candidate answer Let’s think step by step. Lush is a British rock band with four members. P.O.D. is an American rock band with four members. So both Lush and P.O.D. consist of four band members. So the answer is: Yes.
Search query number of members in Lush band + number of members in P.O.D. band
Evidence summary Lush originally consisted of four members: Miki Berenyi, Emma Anderson, Steve Rippon, and Chris Acland. P.O.D. currently consists of four members: Sonny Sandoval, Traa Daniels, Marcos Curiel, and Wuv Bernardo. However, another source mentions only two members of Lush, suggesting a potential inconsistency.
Reflection The evidence confirms P.O.D. has four members, but Lush’s member count is inconsistent across sources. The model failed to resolve the contradiction, leading to a flawed decision.
Refined Query (Iteration 2) Lush band members timeline 1987-1998 + P.O.D. band members lineup history
Evidence summary (Iteration 2) One source claims Lush had five members initially, including Meriel Barham, but another states they had four, with Steve Rippon later replaced by Phil King. P.O.D. remains consistently reported as a four-member band.
Reflection (Iteration 2) The conflicting evidence remains unresolved. Instead of recognizing the uncertainty, the model inaccurately concluded that both bands had four members.
Final decision False (Incorrect)
Explanation Despite P.O.D.’s confirmed four-member structure, the model ignored evidence indicating Lush’s changing member count. It failed to account for the time-specific membership changes, which contradicted the claim that both bands "consist" of four members. This decision highlights TALE’s difficulty in resolving contradictory information.
Table 13: Example of conflicting evidence where TALE failed to resolve contradictions in band member counts.

### v2

@@
 Yet, as these models grow in complexity and capability, achieving timely, affordable, and accurate evaluation of their outputs remains a significant challenge (Xu et al., 2025).
-Consider the question: “What is the tallest building in the world?” The correct answer has changed over time—from the “Willis Tower” to “Taipei 101”, and more recently, to the “Burj Khalifa.” and could be expressed in various valid ways: “Burj Khalifa,” “Khalifa Tower,” “Burj Dubai,” or with its precise height.
-Traditional reference-based evaluation metrics, including Exact Match (EM), F1, BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004), rely on fixed ground-truth references and thus may incorrectly penalize such legitimate variations (Kamalloo et al., 2023; Wang et al., 2023).
-These methods, though efficient, struggle to reflect semantic equivalence, are limited in adaptability to evolving facts.
-On the other hand, human evaluation is generally reliable, but it becomes overly expensive and labor-intensive when applied at scale (Chiang & Lee, 2023; Mañas et al., 2024; Zhu et al., 2023).
-The challenge of evaluating LLMs grows as they act as autonomous agents, handling tasks like web browsing, planning, and information synthesis (Xi et al., 2023; Fu et al., 2023).
+Traditional evaluation methods, including human evaluation and supervised signal-based metrics such as Exact Match (EM), F1, BLEU (Papineni et al., 2002), or ROUGE (Lin, 2004), rely on comparisons to static ground-truth references.
+While convenient and efficient, these methods fall short of capturing the diversity of valid outputs and are often infeasible to scale due to the high cost of human annotations (Chiang & Lee, 2023; Mañas et al., 2024; Zhu et al., 2023).
+The challenge of evaluating LLMs output grows as they act as autonomous agents, handling tasks like web browsing, planning, and information synthesis (Xi et al., 2023; Fu et al., 2023).
 In such dynamic settings, outputs are often unpredictable, context-dependent, and non-deterministic, making it impractical to pre-annotate reference answers for every possible interaction (Li et al., 2024).
@@
 This active evidence-gathering process provides three key advantages: (1) it reduces dependency on the judge’s parameter knowledge, (2) it grounds evaluations in current, verifiable information, and (3) it enables assessment of novel or rapidly evolving topics where parameter knowledge might be outdated or incomplete.
-Our key contributions include:
-  • We introduce TALE, a novel framework that augments LLM-based evaluation with iterative, tool-assisted evidence gathering to overcome the limitations of both reference-based metrics and ungrounded LLM self-judgment.
-  • We implement and evaluate a multi-step reasoning process, enabling LLM judges to actively verify factual claims rather than relying solely on internal knowledge
-  • We demonstrate through extensive experiments that TALE is aligned with reference-based metrics such as F1 and achieves substantial to perfect agreements with human evaluators.
-  • We provide ablation studies that quantify the impact of iterative refinement and query generation capabilities on evaluation performance.
-The results offer insights into the key components that drive TALE’s effectiveness.
+Our key contributions include: (1) We introduce TALE, a novel framework that augments LLM-based evaluation with iterative, tool-assisted evidence gathering to overcome the limitations of both reference-based metrics and ungrounded LLM self-judgment.
+(2) We implement and evaluate a multi-step reasoning process, enabling LLM judges to actively verify factual claims rather than relying solely on internal knowledge.
+(3) We demonstrate through extensive experiments that TALE is aligned with reference-based metrics such as F1 and achieves substantial to perfect agreements with human evaluators.
 2 Related Work
@@
 More closely related to our work, FActScore (Min et al., 2023) decomposes generated text into atomic facts and verifies each against retrieved external knowledge.
+Concurrently, agent-as-a-judge (Zhuge et al., 2024) introduces a tool-augmented judge-agent that walks through an agent’s entire execution trace and labels hundreds of explicit requirements.
+It reaches near-human accuracy while reducing evaluation cost by roughly two orders of magnitude.
 While these approaches leverage external information, they differ fundamentally from TALE in several aspects.
-First, most focus on evaluation through fact extraction and verification rather than holistic response assessment.
-Second, they typically perform single-pass evidence retrieval without the iterative refinement that characterizes TALE.
-Finally, they often lack explicit reasoning about what information to retrieve, instead relying on keyword-based or embedding similarity retrievals.
-Our TALE framework addresses these gaps by implementing an Agent-as-a-judge paradigm (Zhuge et al., 2024) with iterative, adaptive evidence gathering, explicit reflection, and comprehensive reasoning, enabling more robust evaluation in scenarios where static reference answers are unavailable or insufficient.
+Most importantly, previous LLM-as-a-Judge methods output subjective plausibility or preference scores, whereas TALE delivers objective true/false decisions, each accompanied by a rationale grounded in an automatically assembled, evidence-backed checklist.
+Beyond this, earlier approaches score isolated facts rather than the full answer, issue a one-shot evidence query instead of TALE’s iterative retrieve–refine loop, and treat retrieval as a keyword-matching black box rather than a planned, reasoning-driven step.
 3 Methodology
@@
 Figure 1 illustrates the overall process of TALE.
-Figure 1: Given the question, “What is the tallest building in the world?” and the candidate answer, “The tallest building in the world is the Burj Khalifa in Dubai.”, the process begins with an initial query generation based solely on the question to extract relevant information.
-This query is used for evidence gathering through web searches across multiple sources.
-Once the evidence is collected, it is processed by the summary module, which condenses it and extracts key insights.
-The summarized evidence is then passed to the reflector module, which assesses the sufficiency and relevance of the evidence.
-In case of insufficient or conflicting evidence, it triggers a query refinement step to improve the evidence through further iterations.
-After N iterations, the system enters the judgment phase where the judge synthesizes all available evidence and provides a final decision with rationale.
+Figure 1: Given the question “What is the tallest building in the world?” and candidate answer “The tallest building in the world is the Burj Khalifa in Dubai,”, TALE begins with initial query generation from the question alone.
+The query triggers web searches across multiple sources, followed by evidence summarization to extract key insights.
+The reflector module assesses evidence sufficiency and relevance, triggering query refinement if needed.
+After N iterations, the judgment phase synthesizes all evidence to provide a final decision with rationale.
 3.1 Candidate Response Generation
@@
 3.2 LLM-as-a-Judge
-In our framework, an LLM is employed as a judge, denoted by $J$, to assess the correctness of $\hat{y}$ based on external evidence.
-Instead of relying on a fixed reference, the judge acts as an autonomous agent that leverages the aggregated evidence $E$ and reflection $R$, obtained through an iterative process, along with the input question $x$ and the candidate answer $\hat{y}$.
-Formally, the judge produces a binary verdict $v\in\{0,1\}$ (i.e., True or False) and a brief natural language rationale $r_{J}$, such that:
-$(v,r_{J})=J(x,\hat{y},E,R).$
-Here, $J$ is a mapping defined as
-$J:\mathcal{X}\times\mathcal{Y}\times\mathcal{E}\rightarrow\{0,1\}\times\mathbb{R},$
-where $v=1$ indicates that the candidate answer $\hat{y}$ is deemed correct based on the evidence $E$, and $v=0$ otherwise.
-The judge gives precedence to evidence gathered from external sources but may fall back on their pre-trained knowledge when the evidence is incomplete or inconclusive.
+In our framework, an LLM is employed as-a-judge, denoted by $J$, to assess the correctness of $\hat{y}$ based on external evidence.
 3.3 Tool-Augmented Evaluation Process
@@
 The goal is to retrieve background information relevant to the question domain, without assuming the correctness of any particular answer.
-For example, given the question “What is the capital of Switzerland?”, the model may generate a search query like “Capital of Switzerland” or “Switzerland official capital city”.
-This ensures that the evaluation process begins with a neutral, answer-agnostic evidence-gathering step.
 In subsequent iterations ($i>1$), queries are refined using accumulated evidence and reflections stored in the short-term memory: $q_{i}=Q^{\prime}(\mathcal{M})$.
-The purpose is to resolve remaining uncertainty or retrieve more targeted information needed to verify or refute the candidate answer $\hat{y}$.
-For instance, if the candidate’s answer claims “Bern is the capital of Switzerland” but initial evidence suggests there might be a distinction between political and economic capitals, a refined query might be “Is Bern the official capital of Switzerland or is there another capital”.
+The purpose of query refinements is to resolve remaining uncertainty or retrieve more targeted information needed to verify or refute the candidate answer $\hat{y}$.
 Web Search ($S$):
@@
 The search results typically include titles, snippets, and source URLs, providing diverse perspectives from multiple information sources.
-This module enables TALE to access up-to-date and diverse information sources beyond the model’s pre-trained knowledge.
 Summarization ($\Sigma$):
@@
 The resulting summary $E_{i}$ serves as a concise, interpretable knowledge unit that informs reflection.
-For example, a summary might state: “Multiple sources confirm that Bern is the de facto capital and seat of the federal government of Switzerland, though Switzerland does not have an official capital city designated by its constitution.” In TALE, this step is critical for managing context length and ensuring that only high-signal evidence is retained throughout the iterative reasoning process.
 Reflection ($R$):
@@
 Short-Term Memory:
-After each iteration, the tuple $(q_{i},E_{i},R_{i})$—consisting of the generated query, corresponding evidence, and reflection—is appended to the short-term memory buffer:
-$\mathcal{M}\leftarrow\mathcal{M}\cup\{(q_{i},E_{i},R_{i})\}$
+After each iteration, the tuple $(q_{i},E_{i},R_{i})$—consisting of the generated query, corresponding evidence, and reflection—is appended to the short-term memory buffer: $\mathcal{M}\leftarrow\mathcal{M}\cup\{(q_{i},E_{i},R_{i})\}$.
 This memory $\mathcal{M}$ serves as an evolving trace of the evaluation process, accumulating context that informs subsequent query refinement and reasoning.
@@
 Specifically, it includes the aggregated evidence summary and reflections.
-Based on such prior components, $J$ produces a binary verdict $v\in\{0,1\}$ indicating whether $\hat{y}$ is factually correct, along with a natural language rationale $r_{J}$:
-$(v,r_{J})=J(x,\hat{y},E_{\text{total}},R_{\text{total}}).$
+Based on such prior components, $J$ produces a binary verdict $v\in\{0,1\}$ indicating whether $\hat{y}$ is factually correct, along with a natural language rationale $r_{J}$: $(v,r_{J})=J(x,\hat{y},E_{\text{total}},R_{\text{total}})$.
 The overall procedure is summarized in Algorithm 1.
@@
 4 Experimental Setup
-This section details our experimental setup for evaluating TALE across multiple dimensions.
 4.1 Models
@@
 All experiments are conducted with a temperature of 00 to maximize determinism and reliability, as prior work demonstrates that higher temperatures degrade the performance of LLM-based evaluators (Hada et al., 2024).
-For brevity, we refer to these models as Mistral, Llama, Gemini, GPT-3.5, and GPT-4o throughout our analysis.
+For brevity, we refer to these models as Mistral, Gemini, GPT-3.5, and GPT-4o throughout our analysis.
 4.2 Datasets
@@
 We also used FreshQA (Vu et al., 2023) to see TALE’s ability in detecting outdated knowledge.
-Due to computational constraints, we randomly sampled 300 instances from each dataset to ensure balanced representation across question types and difficulty levels.
-Each dataset provides reference answers that serve as ground truth for our reference-based baseline metrics (details in Appendix A.1).
+Due to computational constraints, we randomly sample 300 instances from each dataset, ensuring balanced representation across question types and difficulty levels.
+Each dataset provides reference answers that serve as ground truth for our reference-based baseline metrics (see Appendix A.1).
 4.3 Prompts
@@
 For the TALE framework, we design module-specific prompts that combine role instructions with step-by-step reasoning guidance.
-Each module (query generation, evidence summarization, reflection, refinement, and judgment) uses a one-shot prompt template that demonstrates the expected input-output behavior.
+Each module (query generation, evidence summarization, reflection, and judgment) uses a one-shot prompt template that demonstrates the expected input-output behavior.
 These prompts are crafted to encourage explicit reasoning and evidence-focused analysis without assuming the correctness of any particular answer.
-Complete prompt templates and examples are provided in Appendix A.2.
+See Appendix A.2 for examples.
 4.4 Baselines
@@
 Additionally, we adopt a judge without tool-augmentation (G-Eval) as a baseline, following the approach from Liu et al.
-(2023), where the judge LLM evaluates candidate answers based solely on the question-answer pair without access to external tools or reference answers.
+(2023).
 In this setting, the judge relies entirely on its pre-trained knowledge to determine factual correctness.
-By maintaining the same judge model while removing the evidence retrieval mechanism, this baseline isolates the impact of tool augmentation (details in Appendix A.3).
+By maintaining the same judge model while removing the evidence retrieval mechanism, this baseline effectively isolates the impact of tool augmentation in TALE (details in Appendix A.3).
 Human Evaluation.
@@
 To assess TALE’s performance, we used accuracy as the proportion of instances where the judge’s binary verdict (correct/incorrect) aligns with the ground truth derived from reference answers.
-For AmbigQA and HotpotQA that include human annotations, we calculate Cohen’s Kappa ($\kappa$) and Macro-F1 scores to evaluate TALE’s alignment with human majority votes, with Kappa accounting for agreement beyond chance and Macro-F1 addressing class balance.
-Additionally, we conduct ablation studies to quantify the impact of removing or modifying specific TALE components, using changes in agreement with human judgments as the primary measure.
-For component ablation studies, we measure the change in agreement with human judgments when removing or modifying specific TALE modules.
+For AmbigQA and HotpotQA that include human annotations, we calculate Cohen’s Kappa ($\kappa$) and Macro-F1 scores to evaluate TALE’s alignment with human majority votes.
+Additionally, we conduct ablation performance studies to quantify the impact of removing or modifying specific TALE components, using changes in agreement with human judgments as the primary measure.
 5 Results
-This section briefly presents the results obtained from our experiments. We included additional results and analysis in the Appendix B.
+This section briefly presents the results obtained from our experiments. We included additional results and analysis in the Appendix B and further ablations in the Appendix C.
 By integrating external evidence into pre-trained knowledge, TALE significantly outperforms judges without tools.
@@
 In contrast, TALE consistently achieves the highest Macro F1 across models and tasks.
+TALE works better with more powerful LLMs.
+TALE’s performance improves significantly when using more capable judge models.
+As shown in Tables 2 and 3, GPT-4o consistently outperforms GPT-3.5 and Gemini across both AmbigQA and HotpotQA.
+For instance, GPT-4o within TALE achieves a Cohen’s $\kappa$ of 0.701 on HotpotQA when evaluating GPT-4o as a candidate, compared to 0.540 with GPT-3.5 and 0.679 with Gemini.
+Similarly, GPT-4o reaches a macro F1 of 0.898 on AmbigQA when judging GPT-3.5, higher than GPT-3.5’s 0.784 and Gemini’s 0.830.
 Candidate Task EM F1 Judge w/o Tool (Acc.) TALE (Acc.)
@@
 Table 1: Raw performance of candidate LLMs obtained through different evaluators.
-Evaluation approaches include: (1) Reference-based metrics (EM, F1), (2) Judge w/o Tool (accuracy), and (3) TALE (accuracy).
-TALE works better with more powerful LLMs.
-TALE’s performance improves significantly when using more capable judge models.
-As shown in Tables 2 and 3, GPT-4o consistently outperforms GPT-3.5 and Gemini across both AmbigQA and HotpotQA.
-For instance, GPT-4o within TALE achieves a Cohen’s $\kappa$ of 0.701 on HotpotQA when evaluating GPT-4o as a candidate, compared to 0.540 with GPT-3.5 and 0.679 with Gemini.
-Similarly, GPT-4o reaches a macro F1 of 0.898 on AmbigQA when judging GPT-3.5, higher than GPT-3.5’s 0.784 and Gemini’s 0.830.
 TALE can detect untruthful facts and outdated knowledge.
@@
 Using this information, TALE correctly rejected the candidate’s response, providing a clear rationale grounded in the evidence.
-To evaluate TALE’s ability to detect outdated knowledge, we included FreshQA (Vu et al., 2023), Based on a random sample of 300 questions, we found that TALE consistently identified outdated information in candidate model responses.
+To evaluate TALE’s ability to detect outdated knowledge, we included FreshQA (Vu et al., 2023).
+Based on a random sample of 300 questions, we found that TALE consistently identified outdated information in candidate model responses.
 For instance, when asked “Where is EMNLP this year?”, candidate models often provided outdated responses based on their training data.
@@
 TALE’s reflection process enables it to detect these inconsistencies (see Table 8).
+5.1 Error analysis
+To better understand the limitations of TALE in evaluating candidate responses, we conducted a manual error analysis.
+We randomly sampled 100 evaluation cases from the AmbigQA and HotpotQA datasets, focusing on instances where TALE disagreed with human annotators.
+We categorized the errors into the following categories: 1) Contextual misunderstanding: TALE generates inaccurate or incomplete queries when it misinterprets the intent of the candidate’s question.
+This is particularly evident in AmbigQA, where questions are often intentionally ambiguous or lack sufficient context, leading to the retrieval of irrelevant or contradictory evidence.
+2) Incomplete evidence: TALE fails when the retrieved evidence is insufficient or lacks relevant information, specifically for recent events with limited online coverage.
+3) Reasoning error: Despite accurate evidence, the judge model misinterprets the information or applies flawed reasoning.
+4) Hallucination: In cases where evidence is ambiguous or inconclusive, TALE relies on its pre-trained knowledge, resulting in hallucinated rationales.
+5) Conflicting evidence: In some cases, TALE encounters conflicting evidence across multiple search iterations.
+While the framework is designed to iteratively refine its understanding and gather reliable information, judges sometimes over-rely on earlier sources or fail to appropriately weigh the credibility of conflicting information (details in Appendix C.3).
 Candidate Task EM F1 Judge w/o Tool TALE
@@
 Table 3: Macro F1 between the human majority vote and evaluators. F1 scores are converted to binary using a $\tau=0.5$.
-5.1 Error analysis
-To better understand the limitations of TALE in evaluating candidate responses, we conducted a manual error analysis.
-We randomly sampled 100 evaluation cases from the AmbigQA and HotpotQA datasets, focusing on instances where TALE disagreed with human annotators.
-We categorized the errors into the following categories: 1) Contextual misunderstanding: TALE generates inaccurate or incomplete queries when it misinterprets the intent of the candidate’s question.
-This is particularly evident in AmbigQA, where questions are often intentionally ambiguous or lack sufficient context, leading to the retrieval of irrelevant or contradictory evidence.
-2) Incomplete evidence: TALE fails when the retrieved evidence is insufficient or lacks relevant information, specifically for recent events with limited online coverage.
-3) Reasoning error: Despite accurate evidence, the judge model misinterprets the information or applies flawed reasoning.
-4) Hallucination: In cases where evidence is ambiguous or inconclusive, TALE relies on its pre-trained knowledge, resulting in hallucinated rationales.
-5) Conflicting evidence: In some cases, TALE encounters conflicting evidence across multiple search iterations.
-While the framework is designed to iteratively refine its understanding and gather reliable information, judges sometimes over-rely on earlier sources or fail to appropriately weigh the credibility of conflicting information (details in Appendix B.5).
 5.2 Effect of iterations
-We conducted an ablation study to evaluate the effect of the number of iterations in the TALE framework.
+We conducted an ablation study to evaluate the effect of the number of iterations in TALE.
 Figure 2 shows that increasing the number of iterations generally improves the judge performance across candidate models on both AmbigQA and HotpotQA.
@@
 HotpotQA 0.7993 0.8457
-Table 4: Comparison of Macro F1 scores between Judge w/o Query (GPT-4o as-as-judge) and TALE (GPT-4o as-as-judge).
+Table 4: Macro F1 scores between Judge w/o Query and TALE.
 Note that directly using the input question without generating refined queries is still considered a form of querying, but for clarity, we refer to this setting as w/o Query.
+5.4 Cost analysis
+With three iterations, TALE issues three Serper queries and consumes 4,410 input plus 989 output LLM tokens per answer.
+Using GPT-4o-mini pricing (\$0.15 per $10^{6}$ input tokens and \$0.60 per $10^{6}$ output tokens) together with Serper pricing (\$0.30 per $10^{3}$ queries), this amounts to \$0.00126 (LLM) + \$0.00090 (search) = \$0.00216 per item.
+Evaluating all 300 AmbigQA instances therefore costs \$0.65 and, at 27 s per instance, completes in about 2.3 h.
+Three human annotators need 15 h (300 items $\times$ 3 min) and \$225 (at \$15 h-1).
+Thus TALE is ~6.7$\times$ faster and ~350$\times$ cheaper while retaining human-level accuracy ($\kappa\approx 0.91$).
 6 Conclusion
-We presented TALE, a novel framework for evaluating LLMs.
+We presented TALE, a framework for evaluating LLMs.
 TALE integrates external evidence via an iterative process.
-Through extensive experiments, we found that TALE achieves substantial to perfect agreement with human annotators.
+Through experiments, we found that TALE achieves substantial to perfect agreement with human annotators.
 TALE is also aligned with reference-based metrics such as F1.
@@
 Addressing these biases and ensuring robust evidence selection will be essential for broader deployment.
-Acknowledgments
-We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI), and Research Nova Scotia.
-Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada.
 Appendix A Experimental detail
@@
 We employ a template-based prompting strategy for both response generation and evaluation.
-For candidate models, we utilize few-shot Chain-of-Thought (CoT) prompts, incorporating 6 examples per dataset to encourage detailed, reasoning-driven, and structured responses (see Figure 3).
+For candidate models, we utilize few-shot Chain-of-Thought (CoT) prompts (Gou et al., 2024), incorporating 6 examples per dataset to encourage detailed, reasoning-driven, and structured responses (see Figure 3).
 Candidate Answer Generation Instructions: Answer the following questions.
@@
 So the answer is: Yes.
-Figure 3: Examples of few-shot Chain-of-Thought (CoT) prompts for candidate answer generation.
-TALE prompts candidate LLMs using few-shot Chain-of-Thought (CoT) reasoning to elicit faithful and interpretable outputs across each module.
+Figure 3: Examples of few-shot CoT (Gou et al., 2024) prompts for candidate answer generation.
+TALE prompts candidate LLMs using few-shot CoT reasoning to elicit faithful and interpretable outputs across each module.
 Each module in the evaluation pipeline is guided by a carefully constructed prompt, often in a one-shot format, that combines role-playing instructions with explicit reasoning goals.
@@
 Despite these challenges, Mistral 7B remains a valuable option for resource-constrained environments where efficient evaluation is prioritized over peak accuracy (Badshah & Sajjad, 2024).
+Although Mistral 7B exhibits limitations in complex reasoning and instruction following, it still achieves meaningful alignment with human judgments and surpasses no-tool baselines.
+Since evaluation typically runs offline, users can select a judge model that balances performance and resource constraints.
+Furthermore, TALE’s model-agnostic design ensures that it remains compatible with ongoing improvements in smaller open-source LLMs, enhancing its accessibility and long-term utility.
 Candidate Model Task Accuracy Cohen’s Kappa Macro F1
@@
 While the candidate’s answer concludes correctly, the system highlights the factual inaccuracy in the intermediate reasoning.
-B.5 TALE failure cases
+Appendix C Additional ablations and analysis
+This section presents three extensions to our study.
+First, we compare TALE with three stronger non-iterative baselines to evaluate its relative performance.
+Second, we assess the framework’s robustness to prompt variations.
+Third, we analyze failure cases to better understand the limitations and potential areas for improvement in TALE.
+C.1 Comparison with stronger non-iterative baselines
+We evaluate three non-iterative baselines that exclude TALE’s iterative summarise–reflect–refine process but retain the judgment prompt.
+These methods help disentangle the contributions of tool augmentation, sampling-based self-consistency, and model diversity.
+Across all baselines, we use GPT-4o to generate the candidate answer.
+For the Single-Pass Tool-Augmented Judge and Self-Consistency settings, GPT-4o also serves as the judge model.
+In the Multi-LLM Majority Voting setup, GPT-4o is the candidate model, while three different judge models—GPT-4o, GPT-3.5, and Mistral 7B—independently evaluate the same input to support diverse model reasoning.
+C.1.1 Single-Pass tool-augmented judge
+To isolate the effect of tool access alone, we introduce a baseline that issues exactly one web query and performs no further reasoning or iteration.
+Given a question $q$ and candidate answer $\hat{y}$, the judge LLM formulates a single Serper search, retrieves the top-3 snippets, and immediately produces a true/false verdict with rationale.
+All prompt instructions are identical to those in the full TALE pipeline; the only change is the removal of the summarise–reflect–refine loop.
+This baseline, therefore, quantifies the benefit of a one-shot evidence grab, sitting midway between the Judge without Tool-Augmentation and the full iterative framework.
+As shown in Table 9, on AmbigQA, this naive tool baseline raises Cohen’s $\kappa$ from 0.381 (judge without tools) to 0.658, confirming that one round of external evidence already yields a substantial improvement.
+However, TALE’s pushes $\kappa$ further to 0.914.
+A similar pattern appears on HotpotQA, where $\kappa$ climbs from 0.375 (no tool) to 0.583 (single-pass) and then to 0.701 with TALE.
+C.1.2 Self-consistency judge
+For Self-Consistency (Wang et al., 2023), we let the GPT-4o judge sample 10 independent verdicts at temperature 0.7 and take a simple majority.
+In self-consistency, we excluded an external tool or reference answer, so the model relies exclusively on its parametric knowledge.
+Table 9 shows that self-consistency gives a modest boost over the vanilla judge (no tool), but it still lags far behind the tool-augmented baselines, confirming that external evidence, especially when gathered iteratively is critical for reliable objective judgment.
+C.1.3 Multi-LLM majority voting
+Inspired by PoLL (Verga et al., 2024) and DAFE (Badshah & Sajjad, 2025), we instructed three different LLMs—GPT-4o, GPT-3.5, and Mistral 7B—with the same question and candidate answer, asking each to provide a True/False verdict along with a brief explanation.
+By applying majority voting across the three model outputs, we determined the final answer.
+This ensemble approach leverages the diverse reasoning patterns of different models, increasing overall robustness and reducing individual model biases.
+This approach produces $\kappa$ = 0.609 on AmbigQA and 0.527 on HotpotQA, outperforming self-consistency but remaining below the single-pass tool ($\kappa$ = 0.658/0.583) and far below TALE ($\kappa$ = 0.914/0.701).
+Task No Tool Self-Consistency Multi-LLMs Single-Pass TALE
+AmbigQA 0.381 0.515 0.609 0.658 0.914
+HotpotQA 0.375 0.493 0.527 0.583 0.701
+Table 9: Cohen’s $\kappa$ agreement between LLM-based evaluators and the human majority vote.
+TALE achieves the highest alignment with human judgments across both AmbigQA and HotpotQA.
+Single-Pass denotes tool use without iteration and TALE components (i.e., summarize, reflect, and refine).
+C.2 Prompt variations
+To evaluate TALE’s robustness to prompt design, we re-ran the full framework (with GPT-4o as both candidate and judge) under three prompt variants:
+  • Original (1-shot CoT): Our default configuration, which includes detailed instructions and a single CoT example per module.
+  • Few-shot CoT: An extended version of the prompt containing three CoT examples instead of one.
+  • Simplified (0-shot w/o CoT): Prompt that removes CoT examples for every module.
+As shown in Table 10, the few-shot prompt yields a modest improvement over the original 1-shot setup.
+Removing CoT examples entirely results in a performance drop of at most 3 percentage points in both Cohen’s $\kappa$ and macro-F1.
+Interestingly, the default configuration remains highly stable: re-running TALE with the original prompt reproduced the same $\kappa$ and F1 scores as reported in prior experiments, underscoring its consistency.
+Task Prompt Variant Cohen’s $\kappa$ Macro-F1
+AmbigQA Original (1-shot CoT) 0.914 0.957
+Few-shot CoT 0.920 0.960
+Simplified (0-shot) 0.882 0.932
+HotpotQA Original (1-shot CoT) 0.701 0.850
+Few-shot CoT 0.706 0.853
+Simplified (0-shot) 0.672 0.824
+Table 10: TALE performance under prompt variations. The default (1-shot CoT) is robust, while removing CoT examples yields a small but consistent performance drop.
+C.3 TALE failure cases
 To better understand the limitations of TALE in evaluating candidate responses, we conducted a manual error analysis.
@@
 Table 14: Example of hallucination where TALE falsely identified Sarah Michelle Gellar as a star of Roseanne.
+At each round, TALE’s reflector module takes the aggregated evidence summary and generates a chain-of-thought reflection, pointing out which attributes seem well-supported, which appear contradictory or missing, and where further detail is needed (Section 3.3).
+The next query is generated from that reflection, so contradictions are an explicit signal to search for clarifying evidence.
+We do not hard-code a credibility score; instead, TALE relies on cross-source agreement and iterative follow-up.
+All retrieved snippets and their domains will be released so that future work can plug in credibility weighting without altering the loop.
 Conflicting evidence.
-In some cases, TALE encounters conflicting evidence across multiple search iterations.
-While the framework is designed to iteratively refine its understanding and gather reliable information, it sometimes struggles to appropriately resolve contradictory information.
+At each round, TALE’s reflector module takes the aggregated evidence summary and generates a CoT reflection, pointing out which attributes seem well-supported, which appear contradictory or missing, and where further detail is needed.
+The next query is generated from that reflection, so contradictions are an explicit signal to search for clarifying evidence.
+However, in some cases, TALE encounters conflicting evidence across multiple search iterations.
+It sometimes struggles to appropriately resolve contradictory information.
 Table 15 presents a scenario where the model failed to handle conflicting evidence when assessing the number of band members in Lush and P.O.D.
@@
 Final decision False (Incorrect)
-Explanation Despite P.O.D.’s confirmed four-member structure, the model ignored evidence indicating Lush’s changing member count.
+Table 15: Example of conflicting evidence where TALE failed to resolve contradictions in band member counts.
+Despite P.O.D.’s confirmed four-member structure, the model ignored evidence indicating Lush’s changing member count.
 It failed to account for the time-specific membership changes, which contradicted the claim that both bands "consist" of four members.
 This decision highlights TALE’s difficulty in resolving contradictory information.
-Table 13: Example of conflicting evidence where TALE failed to resolve contradictions in band member counts.
+Topic drift in query refinement.
+To analyze how often TALE’s query refinement step introduces topic drift or irrelevant queries, we examined this within the 100 instances from our error analysis.
+Since our TALE configuration uses up to 3 iterations (1 initial query, up to 2 refined queries), this involved inspecting the query generation process across these stages.
+We found that (see Table 16):
+  • Initial queries, generated directly from the input question, were generally on-topic (sufficient) and relevant to the core question being asked.
+However, due to the ambiguity of some questions, we noticed some off-topic queries in this stage.
+  • Topic drift, while not the most dominant error, did occur during the refinement stages (iterations 2 and 3).
+Out of the 200 refined queries analyzed, only 6 queries in the second iteration and 7 in the third iteration (a total of 13 out of 200, or 6.5%) were judged “insufficient” (indicative of topic drift or clear irrelevance).
+Query Category Initial Query 2nd Query 3rd Query
+Sufficient (On-topic) 77 69 70
+Partially Sufficient 15 25 22
+Insufficient (Off-topic) 8 6 7
+Table 16: Topic drift across 300 queries (100 instances, 3 queries per instance = 300).
+Appendix D Limitations and future work
+Although TALE significantly improves the reliability and accuracy of LLM evaluation in dynamic scenarios, it has several limitations.
+Dependency on judge LLM capabilities.
+TALE’s absolute performance is influenced by the strength of the underlying judge model.
+While our experiments indicate graceful degradation with smaller models like Mistral 7B, the best performance requires stronger models such as GPT-4o.
+Diminishing returns with iterative refinement.
+Our empirical analysis shows that performance gains plateau and sometimes degrade beyond a certain number of iterative evidence-gathering steps due to redundant information and increased strain on the LLM context window.
+Although we currently select an optimal iteration count empirically, future research can develop adaptive stopping criteria, leveraging real-time evidence sufficiency metrics to halt iterations dynamically and efficiently.
+Context window constraints.
+While our short-term memory strategy reduces token input, limitations inherent in current LLM context windows still pose challenges, especially for longer evaluation sequences.
+Future implementations should investigate episodic memory that selectively offloads less relevant information, retrieving only critical snippets required for context-aware judgment, thus more effectively managing context length.
+Source bias and quality control.
+Although TALE’s reflector module actively identifies inconsistencies across multiple evidence sources, our current implementation does not explicitly quantify source credibility.
+To mitigate the impact of source biases and misinformation, future work should prioritize the integration of credibility scoring or weighting mechanisms that dynamically assess evidence reliability.
+Addressing these limitations will enhance TALE’s robustness, scalability, and effectiveness, making it even more valuable for evaluating LLM outputs in diverse, real-world scenarios.
+Appendix E Ethical considerations
+TALE aims to enhance evaluation transparency and factual accuracy by grounding judgments in externally retrieved, verifiable evidence. However, several ethical considerations require attention.
+Reliance on external sources.
+The evidence quality in TALE depends significantly on the retrieved external content, which may introduce unintended biases or inaccuracies.
+Although our iterative refinement explicitly mitigates biases by cross-checking multiple sources, TALE does not currently include explicit credibility scoring.
+Researchers and practitioners should remain cautious about inherent biases in external sources and proactively manage credibility assessments.
+Human oversight and accountability.
+TALE substantially improves automated LLM evaluation accuracy, but human oversight remains critical, especially in high-stakes decision-making contexts.
+It is imperative to recognize the inherent limitations of automated evaluations, maintain human accountability and critical review, particularly when evaluating sensitive or consequential content.
+Computational costs and accessibility.
+TALE’s iterative, tool-augmented process significantly increases computational costs and inference latency compared to simpler evaluation methods, as explicitly shown in our analysis.
+Although this computational overhead is justified by gains in accuracy for offline evaluation scenarios such as benchmarking and auditing, researchers must consider the environmental and resource accessibility implications.
+We encourage future research to optimize computational efficiency and ensure broad accessibility.
+By continuously addressing these ethical dimensions, TALE can support more reliable, transparent, and socially responsible AI evaluation practices.